OpenStudy (anonymous):

What's the difference between least common multiples and least common denominators?

OpenStudy (anonymous):

the least common denominator of two fractions is the least common multiple of the denominators

OpenStudy (anonymous):

So they're the same?

OpenStudy (anonymous):

for example, the least common multiple of 8 and 12 is 24, so the least common denominator of \[\frac{7}{8}\text { and }\frac{5}{12}\] is 24

OpenStudy (anonymous):

How do you find lcm?

OpenStudy (anonymous):

depends on how annoyingly large the numbers are

OpenStudy (anonymous):

I doubt they will be annoyingly large. This is just review, which I am failing at.

OpenStudy (anonymous):

if they are small you can usually guess and check. if they are large, like say 84 and 98 you have to factor each one into products of primes, and then take each prime you see to the highest power you see in any one number

OpenStudy (anonymous):

So I'm subtracting two fractions. I do prime factorization? Then what do I use GCF for?

OpenStudy (anonymous):

so for 8 and 12 i could say \[8=2^3,12=2^2\times 3\implies \text{lcm} =2^3\times 3=24\]

OpenStudy (anonymous):

if you want, and the numbers are not too large, you can always use \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] and cancel at the end

OpenStudy (anonymous):

for example \[\frac{3}{4}-\frac{1}{3}=\frac{3\times 3-4\times 1}{4\times 3}=\frac{9-4}{12}=\frac{5}{12}\]

OpenStudy (anonymous):

7/12 -3/8. So I would prime factorize? And then that would find me the LCM... ?

OpenStudy (anonymous):

we already found the lcm of 8 and 12 is 24

OpenStudy (anonymous):

.... WAT. I know.

OpenStudy (anonymous):

Just confused.

OpenStudy (anonymous):

so you would write \[\frac{7}{12}=\frac{2\times 7}{24}=\frac{14}{24}\] and \[\frac{3}{8}=\frac{3\times 3}{24}=\frac{9}{24}\] and so \[\frac{7}{12}-\frac{3}{8}=\frac{14}{24}-\frac{9}{24}=\frac{5}{24}\]

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

right, i did it the lcm way, not the other way

OpenStudy (anonymous):

Okay, I'm an idiot. Sorry.

OpenStudy (anonymous):

no i am sure you are not an idiot fractions are annoying

OpenStudy (anonymous):

the other way is lcd... I think?

OpenStudy (anonymous):

lcd is the same as lcm, that is you find the lcm of the two denominators and that is your lcd

OpenStudy (anonymous):

Okay, so... I find the lcm of the two denominators by using prime factorization, and once I find that number I kjdfasoudfasuasajsa. Right?

OpenStudy (anonymous):

i guess...